OpenStudy (anonymous):

Help please! B is the midpoint of line segment AC and D is the midpoint of line segment CE. Solve for x, given BD=3x+5 and Ae=4x+20.

OpenStudy (anonymous):



OpenStudy (camper4834):

well i can tell you one thing for sure lets say that the line BD is just B and the line AE is just A i know that B is one half of A

OpenStudy (camper4834):

and we know what B and A are

OpenStudy (anonymous):

How do I fully show my work for this problem? \[(3x+5)=(4x+20)\div2\]

OpenStudy (camper4834):

nevermind!! XD

OpenStudy (anonymous):

Yeah I didn't think I did XD

OpenStudy (camper4834):

idk why i did that!!! XD bad math you are right

OpenStudy (camper4834):

okay so you have it right now you just need to do algebra

OpenStudy (anonymous):

I've tried working it out, and for some reason it just ends up as a mess. I only come here unless its true dire need lol

OpenStudy (camper4834):

\(3x+5=\frac{ 4x+20 }{ 2 } \)

OpenStudy (camper4834):

\(3x+5=2x+10\)

OpenStudy (camper4834):

did you do this step? show me what you did so i can see

OpenStudy (anonymous):

I subtract 5 from both sides, right?

OpenStudy (camper4834):

yup!

OpenStudy (camper4834):

so now we get \(3x=2x+5\)

OpenStudy (anonymous):

Yeah, that's where I'm at.

OpenStudy (camper4834):

do you know that you can subtract 2x from both sides?

OpenStudy (camper4834):

what is 3x - 2x?

OpenStudy (anonymous):

1x, which is actually just x.

OpenStudy (camper4834):

and what is 2x - 2x

OpenStudy (anonymous):

Nothing, they cancel each other out.

OpenStudy (anonymous):

Meaning x=5, right?

OpenStudy (camper4834):

yeah!

OpenStudy (anonymous):

Yhank you so much, I really appropriate that, huge help!

OpenStudy (camper4834):

no problem